Which of the ordered pairs (–10, 3), (5, 0), (–5, –2), (–35, 8)are solutions for the equation x + 5y = 5?
Answer by faceoff57(108)

x + 5y = 5
(-10, 3)
-10+5(3) = 5
-10+15=5
5=5 YES
(5,0)
5+5(0)=5
5=5 YES
(-5, -2)
-5+5(-2)=5
-5-10=5
-15=5 NO
(–35, 8)
-35+5(8)=5
-35+40=5
5=5 YES
(–10, 3), (5, 0), (–35, 8)are solutions for the equation x + 5y = 5